OpenStudy (anonymous):

increasing the volume of a given amount of gas at constant temperature causes the pressure to decrease because

OpenStudy (anonymous):

It is an inverse relationship. PV=PV, temperature is constant. Think about it like this, if you take a balloon that is filled with a certain amount of gas so that it is tight and put that same amount of gas in a bigger balloon (increased volume) the 2nd balloon will not be tight. We put the old amount of gas in a bigger container. It will have less pressure.

OpenStudy (somy):

\(P \times V= n \times R \times T\) \[P= \frac{ n \times R \times T }{ V }\] so if you look at this formula you'll see that if volume is increased - nRT will be divided by bigger number which means the answer for pressure is going to be Less thus increasing volume causes decrease in pressure